find the equation of the parabola determined by the given condition: axis vertical and passing through (0,0), (1,0), and (5,-20)
Answer by ikleyn(52800)

.
find the equation of the parabola determined by the given condition:
axis vertical and passing through (0,0), (1,0), and (5,-20).
~~~~~~~~~~~~~~

From the condition, you see that the quadratic polynomial has zeroes at x= 0 and x= 1.


It means that the quadratic polynomial has this factored form

    y(x) = ax*(x-1),


where "a" is an unknown coefficient.


To find this coefficient, use the third point, which says that  y(5) = -20

    a*5*(5-1) = -20,

or

    20a = -20,

which gives  a = -20/20 = -1.


So, the sough function (polynomial) is  y(x) = -x*(x-1) = -x^2 + x.    ANSWER

Solved.